# When Knockoffs fail: diagnosing and fixing non-exchangeability of Knockoffs

### v1

@@ Abstract
Knockoffs are a popular statistical framework that addresses the challenging problem of conditional variable selection in high-dimensional settings with statistical control. Such statistical control is essential for the reliability of inference. However, knockoff guarantees rely on an exchangeability assumption that is difficult to test in practice, and there is little discussion in the literature on how to deal with unfulfilled hypotheses. This assumption is related to the ability to generate data similar to the observed data. To maintain reliable inference, we introduce a diagnostic tool based on Classifier Two-Sample Tests. Using simulations and real data, we show that violations of this assumption occur in common settings for classical Knockoffs generators, especially when the data have a strong dependence structure. We show that the diagnostic tool correctly detects such behavior. To fix knockoff generation, we propose a nonparametric, computationally-efficient alternative knockoff construction, which is based on constructing a predictor of each variable based on all others. We show that this approach achieves asymptotic exchangeability with the original variables under standard assumptions on the predictive model. We show empirically that the proposed approach restores error control on simulated data.
Controlled variable selection,
Knockoffs,
Classifier Two-Sample tests,
@@
[…]
, , and
@@ 1 Introduction
Controlled variable selection is a fundamental problem encountered in diverse fields where practitioners have to assess the importance of input variables with regards to an outcome of interest. Variable selection aims at choosing variables that should be included in an analysis or model while excluding those that do not substantially contribute to the understanding or prediction of an outcome. Statistically controlled variable selection aims at limiting the proportion of selected variables that are independent of the outcome. This problem is much harder than usual variable selection since guarantees on False Positives have to be provided.
@@
For instance, selecting variables is useful in epidemiology to better understand driving factors in the propagation of diseases (Greenland, 1989). In mathematical finance, choosing an optimal portfolio amounts to solving a variable selection problem (Aït-sahali and Brandt, 2001). In neuroimaging, practitioners are interested in associating cognitive functions to relevant brain regions (Weichwald et al., 2015). In genomics, identifying gene variants that drive disease outcomes is crucial (Visscher et al., 2012). Additionally, False Discoveries have to be controlled to ensure that variable selection is reliable (Benjamini and Hochberg, 1995; Genovese and Wasserman, 2004).
@@
In conditional variable selection, the objective is to assess the value of a particular variable w.r.t. an outcome given a certain set of other variables (König et al., 2021). This approach differs markedly from classical marginal inference, where variables are considered individually in relation to the outcome (Candès et al., 2018). With conditional variable selection, we examine whether a variable remains informative and relevant when considered alongside other variables. This is particularly crucial when dealing with datasets characterized by high levels of correlation among variables. Inference in the presence of correlations is illustrated with a genomics example of Genome-Wide Association Study (GWAS) in Figure 1, where correlation between variables is due to a biological phenomenon called linkage disequilibrium (Uffelmann et al., 2021).
@@
Figure 1: Illustration of conditional variable selection. Given a genomic dataset and kidney stones diagnoses (outcome), we aim at finding variables that add information w.r.t the outcome given all other variables. This problem is challenging since genes are highly correlated locally. Assessing the added information of specific genes given all others is crucial to improve the understanding of disease outcomes. Details about this Manhattan plot are available in Howles et al., 2019.
@@
This problem becomes difficult when there is a high level of correlation between variables on which one aims to perform inference. In machine learning contexts, the difficulty of identifying features that add unique predictive value when combined with others in large-correlation contexts has been well identified (Bickel, Ritov and Tsybakov, 2009; Mandozzi and Bühlmann, 2016; Goeman and Bühlmann, 2007; Javanmard and Montanari, 2014). High correlations are ubiquitous in many application areas, for example when the features of interest are derived from some biological measurements that inherently have complex dependencies. For instance, in brain mapping, practitioners may want to detect brain regions that are relevant for a given cognitive task given the rest of the brain (Weichwald et al., 2015), but have to deal with the strong dependencies observed in these data (Chevalier et al., 2021). In genomics, understanding which genes conditionally affect disease outcomes can lead to more effective disease prevention and intervention strategies (Sesia, Sabatti and Candès, 2019).
@@
Knockoffs (Barber and Candès, 2015; Candès et al., 2018) have emerged as a popular and powerful tool for addressing the problem of conditional variable selection with guarantees on the rate of false detections. This framework consists of guiding variable selection by generating synthetic variables, referred to as knockoffs, that closely mimic the statistical properties of the original variables while not being associated to the outcome of interest (conditionally on their original counterpart). This approach aims to identify genuine associations between variables and outcomes by comparing them with their corresponding knockoffs. An attractive feature of knockoffs is that inference is performed simultaneously for all variables, resulting in an acceptable computational cost. Performing the same type of inference for each variable leads to as many Conditional Randomization Tests (Candès et al., 2018; Nguyen, Thirion and Arlot, 2022) as the number of covariates, which has a high computational cost.
@@
The most popular knockoff generation procedure involves a Gaussian assumption, and thus relies on the knowledge or the accurate estimation of the covariance structure of the covariates, which leads to two potential issues: i) the violation of the Gaussian hypotheses and ii) inaccuracies in the covariance estimation. This worsens with the number of variables $p$ as the problem becomes harder, especially in high-dimensional regimes where the number of samples $n$ is comparatively small. Some alternative Knockoff generation methods avoid the problem of covariance estimation via e.g. Deep Learning (Romano, Sesia and Candès, 2020). However, such methods also suffer from high-dimensional regimes as deep neural networks require massive amounts of data to be properly trained, especially in large feature spaces. Overall, methods that require a large sample size $n$ and a small number of variables $p$ may not be adapted to the Knockoffs framework, which is designed for high-dimensional variable selection. Domain-specific procedures have been developed to tackle some of these issues. For instance, in the field of genomics where variables of interest may be discrete, Knockoffs can be built using Hidden Markov Models (HMM) (Rabiner and Juang, 1986), as demonstrated in the work of Sesia, Sabatti and Candès (2019).
@@
The aim of this work is twofold: First, we propose an effective diagnostic tool that allows practitioners to examine knockoffs along the original data for potential violations. Based on this, we highlight common cases of non-exchangeability when using Gaussian knockoffs and associated error control violations. We describe their consequences, typically a failure to control the false discovery rate at the nominal level. We propose an efficient non-parametric algorithm for constructing knockoffs which comes with theoretical guarantees. We show through simulations that this procedure restores error control in all cases considered.
2 Knockoffs framework and related work
@@
Notation. For an integer $k$, $\llbracket k\rrbracket$ denotes the set $\{1,\dots,k\}$. Equality in distribution is denoted by =$\stackrel{{\scriptstyle d}}{{=}}$. We denote vectors by bold lowercase letters. A vector $\mathbf{x}=(x_{j})_{j\in\llbracket p\rrbracket}$ from which the $j^{th}$ coordinate is removed is denoted by $\mathbf{x}_{-j}$, i.e. $\mathbf{x}\setminus\{x_{j}\}$. Independence between two random vectors $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\perp\mathbf{y}$. For two vectors $\mathbf{x}$ and $\mathbf{\tilde{x}}$ of size $p$ and a subset $S\subset\llbracket p\rrbracket$, $(\mathbf{x},\mathbf{\tilde{x}})_{swap(S)}$ denotes the vector obtained from $(\mathbf{x},\mathbf{\tilde{x}})$ by swapping the entries $x_{j}$ and $\tilde{x}_{j}$ for each $j\in S$. Matrices are denoted by bold uppercase letters, the only exception being the vector of Knockoff statistics that we denote by $\mathbf{W}$ as in Barber and Candès (2015); Candès et al. (2018). For any set $S$, $|S|$ denotes the cardinality of $S$. For a vector $\mathbf{z}=(z_{j})_{1\leq j\leq p}$ and $S\subset\llbracket p\rrbracket$, we denote by $z_{(j:S)}$ (or $z_{(j)}$ when there is no ambiguity) the $j^{th}$ smallest value in the sub-vector $(\mathbf{z}_{s})_{s\in S}$.
Problem setup. The input data are denoted by $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n$ is the number of samples and $p$ the number of variables. The outcome of interest is denoted by $\mathbf{y}\in\mathbb{R}^{n}$. The goal is to select variables that are relevant with regards to the outcome conditionally on all others. Formally, we test simultaneously for all $j\in\llbracket p\rrbracket$:
$H_{0,j}:y\perp x_{j}|\mathbf{x}_{-j}\quad\text{versus}\quad H_{1,j}:y\not\perp x_{j}|\mathbf{x}_{-j}.$
@@
The output of a variable selection method is a rejection set $\hat{S}\subset\llbracket p\rrbracket$ that estimates the true unknown support $\mathcal{H}_{1}=\{j:y\not\perp x_{j}|\mathbf{x}_{-j}\}$. Its complement is the set of true null hypotheses $\mathcal{H}_{0}=\{j:y\perp x_{j}|\mathbf{x}_{-j}\}$. To ensure reliable inference, our aim is to provide a statistical guarantee on the proportion of False Discoveries in $\hat{S}$. The False Discovery Proportion (FDP) and the False Discovery Rate (FDR) Benjamini and Hochberg (1995) are defined as:
@@
Knockoffs. Knockoff inference leverages noisy duplicates known as knockoff variables which serve as controls in the variable selection process. A key challenge in this method is to ensure that these knockoff variables maintain the same correlation structure as the original variables, while being conditionally independent of the outcome. This is essential to enable meaningful comparisons between the original variables and their knockoff counterparts, thereby identifying variables that provide relevant information regarding the outcome. Knockoff variables are defined as follows:
Definition 1 (Model-X Knockoffs, Candès et al., 2018).
For the family of random variables $\mathbf{x}=(x_{1},\ldots,x_{p})$, Knockoffs are a new family of random variables $\mathbf{\tilde{x}}=(\tilde{x}_{1},\ldots,\tilde{x}_{p})$ satisfying:
@@
    for any $S\subset\llbracket p\rrbracket$, $(\mathbf{x},\mathbf{\tilde{x}})_{swap(S)}\stackrel{{\scriptstyle d}}{{=}}(\mathbf{x},\mathbf{\tilde{x}})$
@@
    $\mathbf{\tilde{x}}\perp\mathbf{y}|\mathbf{x}$.
Exchangeability is the first property of this definition. For any set of swapped variables, the joint distribution of real and knockoff variables must remain identical to the original one. Building Knockoff variables that satisfy this property is a difficult problem in itself – we discuss existing work on Knockoff generation and propose a novel approach in Section 3.
Statistical inference on Knockoff variables. Once valid knockoffs have been created, conditional variable selection can be performed. To distinguish variables that are substantially more important than their corresponding Knockoffs, a machine learning model is employed to generate importance scores for each variable and its respective knockoff. This step enables the identification of variables that offer valuable insights into the outcome, as they exhibit significant disparities in importance compared to their knockoffs. Quantitatively, this is done by computing Knockoff statistics $\mathbf{W}=(W_{1},\ldots,W_{p})$ that are defined as follows.
Definition 2 (Knockoff Statistic, Candès et al., 2018).
A knockoff statistic $\mathbf{W}=(W_{1},\ldots,W_{p})$ is a measure of feature importance that satisfies:
@@
    $\mathbf{W}$ depends only on $\mathbf{X},\mathbf{\tilde{X}}$ and $\mathbf{y}$: $\mathbf{W}=g(\mathbf{X},\mathbf{\tilde{X}},\mathbf{y})$.
@@
The most commonly used Knockoff statistic is the Lasso-coefficient difference (LCD) (Weinstein et al., 2020). This statistic is obtained by fitting a Lasso estimator (Tibshirani, 1996) on $[\mathbf{X},\mathbf{\tilde{X}}]\in\mathbb{R}^{n\times 2p}$, which yields $\widehat{\boldsymbol{\beta}}\in\mathbb{R}^{2p}$. Then, the Knockoff statistic can be computed using $\widehat{\boldsymbol{\beta}}$:
@@
This definition of $T_{q}$ ensures that the $\operatorname{FDR}$ is controlled at level $q$ (Candès et al., 2018). Alternatively, inference can be performed using $\pi$-statistics, which quantify the evidence against each variable:
@@
As noted by Nguyen et al. (2020), the vanilla Knockoffs procedure of Candès et al., 2018 is equivalent to using the Benjamini and Hochberg (1995) procedure at level $q$ on the vector of $\pi$-statistics $(\pi_{j})_{j\in\llbracket p\rrbracket}$. The complete procedure is summarized in Figure 2.
Figure 2: Conditional variable selection via Knockoff variables. Knockoff variables are noisy copies of the original variables that are used as controls for variable selection (Panel a). Conditionally on the real variables, knockoffs are independent of the outcome $y$. Importance scores for each variable are computed using the sensitivity estimate from machine learning model – typically, the corresponding coefficient of a Lasso fit. The Knockoff statistic is defined as the difference of modulus between the score associated to the real variable and its Knockoff counterpart (Panel b). Then, $\pi$-statistics are computed to rank variables and to perform inference (Panel c).
Randomness. Beyond the validity of Knockoff variables, another notable challenge is the random nature of the knockoffs generation process. This randomness introduces instability in variable selection outcomes – this is highly undesirable in applications and hampers reproducibility.
Much work has been done recently to tackle this issue, including using stability selection criteria (Ren, Wei and Candès, 2021), aggregating multiple $\pi$-statistics (Nguyen et al., 2020; Blain et al., 2023) and using aggregated Knockoffs e-values (Ren and Barber, 2022). The gist of these approaches is to draw multiple sets of Knockoffs and aggregate the resulting statistics or rejection sets in order to robustify the variable selection process. Such methods offer ways to provide more stable and reliable selection sets, thus mitigating the randomness associated with the knockoff generation process. These advancements enhance the credibility and reproducibility of conditional variable selection using knockoffs.
3 Knockoffs generation methods

### v2

@@ Abstract
Knockoffs are a popular statistical framework that addresses the challenging problem of conditional variable selection in high-dimensional settings with statistical control. Such statistical control is essential for the reliability of inference. However, knockoff guarantees rely on an exchangeability assumption that is difficult to test in practice, and there is little discussion in the literature on how to deal with unfulfilled hypotheses. This assumption is related to the ability to generate data similar to the observed data. To maintain reliable inference, we introduce a diagnostic tool based on Classifier Two-Sample Tests. Using simulations and real data, we show that violations of this assumption occur in common settings for classical knockoff generators, especially when the data have a strong dependence structure. We show that the diagnostic tool correctly detects such behavior. We show that an alternative knockoff construction, based on constructing a predictor of each variable based on all others, solves the issue. We also propose a computationally-efficient variant of this algorithm and show empirically that this approach restores error control on simulated data and semi-simulated experiments based on neuroimaging data.
@@
Controlled variable selection, Knockoffs, Classifier Two-Sample tests
@@ 1 Introduction
Controlled variable selection is a fundamental problem encountered in diverse fields where practitioners have to assess the importance of input variables with regards to an outcome of interest. Variable selection aims at choosing variables that should be included in an analysis or model while excluding those that do not substantially contribute to the understanding or prediction of an outcome. Statistically controlled variable selection aims at limiting the proportion of selected variables that are independent of the outcome. This problem is much harder than usual variable selection since guarantees on False Positives have to be provided. For instance, selecting variables is useful in epidemiology to better understand driving factors in the propagation of diseases (Greenland,, 1989). In mathematical finance, choosing an optimal portfolio amounts to solving a variable selection problem (Aït-sahali and Brandt,, 2001). In neuroimaging, practitioners are interested in associating cognitive functions to relevant brain regions (Weichwald et al.,, 2015). In genomics, identifying gene variants that drive disease outcomes is crucial (Visscher et al.,, 2012). Additionally, False Discoveries have to be controlled to ensure that variable selection is reliable (Benjamini and Hochberg,, 1995; Genovese and Wasserman,, 2004).
@@
In conditional variable selection, the objective is to assess the value of a particular variable w.r.t. an outcome given a particular set of other variables (König et al.,, 2021). This approach differs markedly from classical marginal inference, where variables are considered individually in relation to the outcome (Candès et al.,, 2018). With conditional variable selection, we examine whether a variable remains informative and relevant when considered alongside other variables. This is particularly crucial when dealing with datasets characterized by high levels of correlation among variables. Inference in the presence of correlations is illustrated with a genomics example of Genome-Wide Association Study (GWAS) in Figure 1, where correlation between variables is due to a biological phenomenon called linkage disequilibrium (Uffelmann et al.,, 2021).
@@
Figure 1: Illustration of conditional variable selection. Given a genomic dataset and kidney stones diagnoses (outcome), we aim at finding variables that add information w.r.t the outcome given all other variables. This problem is challenging since genes are highly correlated locally. Assessing the added information of specific genes given all others is crucial to improve the understanding of disease outcomes. Details about this Manhattan plot are available in Howles et al.,, 2019.
@@
This problem becomes difficult when there is a high level of correlation between variables on which one aims to perform inference. In machine learning contexts, the difficulty of identifying features that add unique predictive value when combined with others in large-correlation contexts has been well identified (Bickel et al.,, 2009; Mandozzi and Bühlmann,, 2016; Goeman and Bühlmann,, 2007; Javanmard and Montanari,, 2014). High correlations are ubiquitous in many application areas, for example when the features of interest are derived from some biological measurements that inherently have complex dependencies. For instance, in brain mapping, practitioners may want to detect brain regions that are relevant for a given cognitive task given the rest of the brain (Weichwald et al.,, 2015), but have to deal with the strong dependencies observed in these data (Chevalier et al.,, 2021). In genomics, understanding which genes conditionally affect disease outcomes can lead to more effective disease prevention and intervention strategies (Sesia et al.,, 2019).
@@
Knockoffs (Barber and Candès,, 2015; Candès et al.,, 2018) have emerged as a popular and powerful tool for addressing the problem of conditional variable selection with guarantees on the rate of false detections. This framework consists of guiding variable selection by generating synthetic variables, referred to as knockoffs, that closely mimic the statistical properties of the original variables while not being associated with the outcome of interest (conditionally on their original counterpart). This approach aims to identify genuine associations between variables and outcomes by comparing them with their corresponding knockoffs. An attractive feature of knockoffs is that inference is performed simultaneously for all variables, resulting in an acceptable computational cost. Performing the same type of inference for each variable leads to as many Conditional Randomization Tests (Candès et al.,, 2018; Nguyen et al.,, 2022) as the number of covariates, which has a high computational cost.
@@
The most popular knockoff generation procedure involves a Gaussian assumption, and thus relies on the knowledge or the accurate estimation of the covariance structure of the covariates, which leads to two potential issues: i) the violation of the Gaussian hypothesis and ii) inaccuracies in the covariance estimation. This worsens with the number of variables $p$ as the problem becomes harder, especially in high-dimensional regimes where the number of samples $n$ is comparatively small.
@@
Some alternative knockoff generation methods avoid the problem of covariance estimation via e.g. Deep Learning (Romano et al.,, 2020). However, such methods also suffer from high-dimensional regimes as deep neural networks require massive amounts of data to be properly trained, especially in large feature spaces. Overall, methods that require a large sample size $n$ and a small number of variables $p$ may not be adapted to the knockoffs framework, which is designed for high-dimensional variable selection. Domain-specific procedures have been developed to tackle some of these issues. In particular, motivated by applications to GWAS in genomics, Monte-Carlo based knockoff generation methods have been proposed for discrete distributions (Sesia et al.,, 2019; Bates et al.,, 2021). In the continuous setting, the latter is equivalent to Gaussian knockoffs.
@@
This paper is organized as follows. In Sections 2 and 3, we describe the knockoff framework and review knockoff generation methods. In particular, an efficient nonparametric algorithm for constructing knockoff variables is proposed in Section 3.3. In Section 4, we introduce an effective diagnostic tool that allows practitioners to examine knockoffs along the original data for potential violations. In Section 5 this tool is implemented to illustrate through simulations and application to semi-simulated fMRI data common cases of non-exchangeability when using Gaussian knockoffs and associated error control violations. We also show that the proposed nonparametric knockoff generation procedure restores error control in all the cases considered.
2 Knockoff framework
@@
Notation. For an integer $k$, $\llbracket k\rrbracket$ denotes the set $\{1,\dots,k\}$. Equality in distribution is denoted by =$\stackrel{{\scriptstyle d}}{{=}}$. A vector $X=(X_{j})_{j\in\llbracket p\rrbracket}$ from which the $j^{th}$ coordinate is removed is denoted by $X_{-j}$, that is, $X\setminus\{X_{j}\}$. The independence between two random vectors $X$ and $Y$ is denoted by $X\perp Y$. For two vectors $X$ and $\tilde{X}$ of size $p$ and a subset $S\subset\llbracket p\rrbracket$, $(X,\tilde{X})_{swap(S)}$ denotes the vector obtained from $(X,\tilde{X})$ by swapping the entries $X_{j}$ and $\tilde{X}_{j}$ for each $j\in S$. Matrices are denoted by bold letters. $\mathbf{X}_{j}$ (resp. $\mathbf{X}^{i}$) stands for the $j$-th column (resp. $i$-th row) of matrix $\mathbf{X}$. For any set $S$, $|S|$ denotes the cardinality of $S$. Regressors will be taken from $\mathcal{F}$, which denotes a class of measurable functions. Problem setup. The input data are denoted by $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n$ is the number of samples and $p$ the number of variables. The outcome of interest is denoted by $\mathbf{Y}\in\mathbb{R}^{n}$. The goal is to select variables that are relevant with respect to the outcome conditionally on all others. Formally, we test simultaneously for all $j\in\llbracket p\rrbracket$:
$H_{0,j}:Y\perp X_{j}|X_{-j}\quad\text{versus}\quad H_{1,j}:Y\not\perp X_{j}|X_{-j}.$
@@
The output of a variable selection method is a rejection set $\hat{S}\subset\llbracket p\rrbracket$ that estimates the true unknown support $\mathcal{H}_{1}=\{j:Y\not\perp X_{j}\mid X_{-j}\}$. Its complement is the set of true null hypotheses $\mathcal{H}_{0}=\{j:Y\perp X_{j}\mid X_{-j}\}$. To ensure reliable inference, our aim is to provide a statistical guarantee on the proportion of False Discoveries in $\hat{S}$. The False Discovery Proportion (FDP) and the False Discovery Rate (FDR) Benjamini and Hochberg, (1995) are defined as:
@@
Knockoffs. Knockoff inference leverages noisy duplicates known as knockoff variables, which serve as controls in the variable selection process. A key challenge in this method is to ensure that these knockoff variables maintain the same correlation structure as the original variables while being conditionally independent of the outcome. This is essential to allow meaningful comparisons between the original variables and their knockoff counterparts, thereby identifying variables that provide relevant information regarding the outcome. Knockoff variables are defined as follows:
Definition 1 (Model-X knockoffs, Candès et al.,, 2018)
For the family of random variables $X=(X_{1},\ldots,X_{p})$, knockoffs are a new family of random variables $\tilde{X}=(\tilde{X}_{1},\ldots,\tilde{X}_{p})$ satisfying:
@@
    for any $S\subset\llbracket p\rrbracket$, $(X,\tilde{X})_{swap(S)}\stackrel{{\scriptstyle d}}{{=}}(X,\tilde{X})$
@@
    $\tilde{X}\perp Y|X$.
Exchangeability is the first property of the above definition. For any set of swapped variables, the joint distribution of real and knockoff variables must remain identical to the original one. Building knockoff variables that satisfy this property is a difficult problem in itself – we discuss existing work on knockoff generation in Section 3. The second property is generally achieved by building $\tilde{X}$ from $X$ only.
Statistical inference on knockoff variables. Once valid knockoffs have been created, conditional variable selection can be performed. To distinguish variables that are substantially more important than their corresponding knockoffs, a machine learning model is employed to generate importance scores for each variable and its respective knockoff. This step enables the identification of variables that offer valuable insights into the outcome, as they exhibit significant disparities in importance compared to their knockoffs. Quantitatively, this is done by computing knockoff statistics $W=(W_{1},\ldots,W_{p})$ that are defined as follows.
Definition 2 (Knockoff statistic, Candès et al.,, 2018)
A knockoff statistic $W=(W_{1},\ldots,W_{p})$ is a measure of feature importance that satisfies:
@@
    $W$ depends only on $\mathbf{X},\mathbf{\tilde{X}}$ and $\mathbf{Y}$: $W=g(\mathbf{X},\mathbf{\tilde{X}},\mathbf{Y})$.
@@
The most commonly used knockoff statistic is the Lasso-coefficient difference (LCD) (Weinstein et al.,, 2020). This statistic is obtained by fitting a Lasso estimator (Tibshirani,, 1996) on $[\mathbf{X},\mathbf{\tilde{X}}]\in\mathbb{R}^{n\times 2p}$, which yields $\widehat{\beta}\in\mathbb{R}^{2p}$. Then, the knockoff statistic can be computed using $\widehat{\beta}$:
@@
This definition of $T_{q}$ ensures that the $\operatorname{FDR}$ is controlled at level $q$ (Candès et al.,, 2018). Alternatively, inference can be performed using $\pi$-statistics, which quantify the evidence against each variable:
@@
As noted by Nguyen et al., (2020), the vanilla knockoffs procedure of Candès et al.,, 2018 is equivalent to using the Benjamini and Hochberg, (1995) procedure at level $q$ on the vector of $\pi$-statistics $(\pi_{j})_{j\in\llbracket p\rrbracket}$. The complete procedure is summarized in Figure 2.
Figure 2: Conditional variable selection via knockoff variables. Knockoff variables are noisy copies of the original variables that are used as controls for variable selection (Panel a). Conditionally on the real variables, knockoffs are independent of the outcome $y$. Importance scores for each variable are computed using the sensitivity estimate from machine learning model – typically, the corresponding coefficient of a Lasso fit. The knockoff statistic is defined as the difference of modulus between the score associated with the real variable and its knockoff counterpart (Panel b). Then, $\pi$-statistics are computed to rank variables and to perform inference (Panel c).
3 Knockoff generation methods